# LSHBloom: Internet-Scale Text Deduplication

### v3

@@ 3. Related Work
Having provided background on efficient document similarity measures, we briefly review related work on deduplication methods employed by groups developing LLMs. LLM developers use various methods, many ad hoc, to deduplicate the large quantities of text content used for LLM training regardless of whether the text comes from a traditional source or synthetic data. Many leading LLMs do not discuss the deduplication and curation methods they use, such as Llama-2 (touvron_llama_2023-1), Claude, Gemini (gemini_team_gemini_2024; gemini_team_gemini_2024-1), and GPT versions 3 and later (openai_gpt-4_2024). Others, such as Palm2 (anil2023palm) and Gemma (gemma_team_gemma_2024; gemini_team_gemini_2024-1), provide only limited details. Others, including several leading models, provide extensive details about how data for their model was deduplicated including Olmo’s Dolma dataset (soldaini_dolma_2024) as well as its successor DataComp-LM (li2024datacomp), Llama 1 (touvron_llama_2023), Llama 3 (dubey_llama_2024), and OPT (zhang_opt_2022). For example, Llama 3 (dubey_llama_2024) was trained on a dataset that was deduplicated by using MinHash over the entire dataset, line-level deduplication using CCNet for lines replicated over six times in each bucket of text based on coarse categorization (e.g., math reasoning, geometry/trigonometry), and URL deduplication for web data. Phi uses an approach described by gunasekar_textbooks_2023 and elaborated on in their Phi 2.0 report (javaheripi_surprising_2023) that starts from deduplicated code and text, but they do not discuss deduplicating after generation. C4 (t5_c4_raffel2020) applies only heuristic filtering for data quality (e.g., removing lines without punctuation or fewer than three words). RedPajama (together2023redpajama) uses both exact matching via Bloom filters and fuzzy deduplication via MinHashLSH. Datasets such as the Pile (gao2020pile800gbdatasetdiverse) and more recent releases such as FineWeb2 (penedo2025fineweb2pipelinescale) apply MinHashLSH for fuzzy deduplication.
The most common methods used across these model sources include metadata-based approaches, MinHashLSH, CCNet, and n-gram-based methods. Metadata-based approaches (e.g., exact URL matching, DOIs) are by far the simplest approach; however, they are applicable only in cases in which a reliable source of metadata exists, such as those provided by a single publisher. Unfortunately, for our target datasets this metadata is often unavailable or is otherwise untrustworthy. This can occur, for example, when: 1) a document appears in two or more collections with different metadata schemes (e.g., ArXiV vs. ACM, different GitHub repositories); 2) sections/paragraphs of text are copied, as in extended journal versions of computer science papers; or 3) metadata is not curated (e.g., an ArXiV entry may not be updated to reflect a subsequent publication).
@@
CCNet (wenzek2020ccnet) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content. CCNet performs dataset sharding, deduplication, language identification, and quality filtering. CCNet’s deduplication technique begins by preprocessing text so that all characters are lowercase and eliminates special Unicode characters. CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates. CCNet can optionally compare to a single document, a subset of documents, or all documents.
@@
Dolma (soldaini_dolma_2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma). An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter. If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate. By default, $T$ = 1.0 in their configuration.
@@
DataComp-LM (li2024datacomp) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter. They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer (ye2023uniseg). For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams. If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document. If the proportion of duplicated n-grams in the entire document exceeds the user-defined threshold, then the document is removed from the dataset. They employ the same threshold parameters for paragraph-level and document-level deduplication.
A table comparing various methods for data deduplication in contemporary LLM pre-training pipelines can be found in Appendix A.
4. The LSHBloom Method
4.1. Overview
@@
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of bender2022optimal.
@@ 4.2. Insertion
Given a Jaccard similarity threshold $T\in$ [0, 1], MinHashLSH traditionally groups signatures into a series of $b$ bands of size $r$ such that an exact match in one band for two columns of the MinHash signature matrix denotes a duplicate document. That is, with high probability, the two documents with a matching band have Jaccard similarity greater than or equal to the threshold. To avoid storing the signatures in full, as is required by MinHashLSH, we assign each band of the minhash signature matrix its own Bloom filter. Thus we have in total $b$ Bloom filters, and a collision in any one indicates a duplicate document.
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (sha1)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values. Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (carter1977universal):
@@ 4.3. Querying
Querying a document against the index is a similar process to insertion. We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in subsection 4.2. If there is a collision for any band of signatures, we mark that document as a duplicate. This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters. However, the false positive probability is configurable and can be set to be arbitrarily low.
@@ 4.4. Error Rate
In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate. This is also the case in LSHBloom. However, we make use of Bloom filters, which incur additional false positive overhead. Here, we analyze the error rate of LSHBloom. A false positive is a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom, and a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
Given a Minhash signature matrix with $b$ bands with $r$ rows each corresponding to $b$ Bloom filters with their individual false positive rates set to $p$, the net false positive overhead of using Bloom filters will be $p\textsubscript{effective}=1-(1-p)^{b}$. This is because a single collision in any band constitutes a duplicate, and false positive rates are independent for each Bloom filter. Therefore, the probability that no Bloom filter reports a false positive is $(1-p)^{b}$, and the overall false positive rate overhead introduced by using Bloom filters is the complement of this value.
The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows. leskovec2014finding determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
@@
zhu_lshensemble demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
@@
A Bloom filter will not yield false negatives, but may yield false positives. If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result. If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$. Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.2) for insertion into the Bloom filter. Therefore, the overall false positive rate for LSHBloom is given by:
(3) $FP_{bloom}=FP_{lsh}+(1-FP_{lsh})*(p_{effective}+b/N)$
@@
Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible. For 32-bit hashvalues (as is the default in packages such as datasketch) $N=2^{32}$, so the probability of collision for bands prior to insertion into the Bloom filter is also quite small. Additionally, because the size of the Bloom filter index scales with the number of items inserted and is not influenced by the size of the integers inserted into it, one can increase the size of the hashvalues (to say, a 64 or 128-bit integer) without increasing the size of the Bloom filter index. This is not true of the traditional MinHashLSH index which must store these hashvalues for comparison.
@@
Meanwhile, for each false negative that MinHashLSH produces, $p_{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the bloom filter index and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion. That is,
(4) $FN_{bloom}=(1-(p_{effective}+b/N))*FN_{lsh}$
@@ 4.5. Advantages
We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams. Since many practical scenarios may admit an extremely large universe of such n-grams (e.g., if we allow our set of documents to include a variety of encodings, languages, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the size of MinHash hashvalues (to 64 or 128-bits or more) or increasing the accuracy of MinHash estimates of the Jaccard similarity by increasing the number of MinHash permutations. A scalable deduplication algorithm should be amenable to such considerations.
@@
LSHBloom allows for the space-efficient representation of extremely large deduplication indices. This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with a fixed number of bits. This means that increasing the size of hashvalues does not change the index size. Furthermore, using more MinHash permutations only increases the index size linearly with the number of bands that are required for the MinHash signature matrix–which is far fewer than the number of permutations. If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (bender2022optimal).
@@
As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing. For these parameter settings, MinHashLSH creates nine bands (using the method described in (zhu_lshensemble) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index. For these settings, setting a conservative ppeffective = 1e-10 for a dataset of 10B documents would only require 590 GB of space. Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex—exemplifying the scalability of our technique.
While the use of Bloom filters to satisfy our memory and runtime constraints introduces additional false positive errors, it also provides another level of configuration. This is another advantage of our technique: we can observe massive gains in space while only introducing a single additional hyperparameter, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
@@
The additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter. Therefore, there is a trade-off between memory usage and correctness that is mediated by the size of each filter. However, even for very large datasets of, say, 100M documents, this does not prove to be an issue as we can set an effective false positive probability overhead as low as 1e-15 and only incur a space cost of 8.5 GB for our index compared to 460 GB for the equivalent LSHIndex.

### v4

@@ 3. Related Work
The task of streaming text deduplication lies at the intersection of the literature on Set-Similarity joins, Approximate Membership Queries, and Streaming Entity Resolution. While traditional approaches prioritize fidelity at the cost of computational resources, LSHBloom targets the Pareto-optimal frontier of high-dimensional streaming deduplication—offering state-of-the-art deduplication fidelity while minimizing resource usage.
3.1. Similarity Joins
Traditional techniques for performing a Set-Similarity Join attempt to identify all pairs of records such that their set-similarity (as measured by the Jaccard, Dice, or Cosine metrics) exceeds a given threshold. Mann et al. (2016) perform a comprehensive evaluation of techniques for performing the Set-Similarity Join (e.g., AllPairs, PPJoin). While employing certain filter optimizations can improve upon the quadratic cost of computing exact set-similarity metrics, ultimately, these techniques are not appropriate for extremely high-dimensional entities such as documents. Techniques such as SemDedup (Abbas et al., 2023) aim to deduplicate text entries based on their overlap in terms of semantic content rather than using purely textual overlap to compute similarity; however, due to the use of language models, this technique incurs significant computational costs. Metadata-based approaches (e.g., exact URL matching, matching DOIs), including blocking using common field values (Christen, 2012), are a simple and efficient alternative, but are applicable only when a reliable source of metadata exists. Unfortunately, for our target datasets metadata is often unavailable or is untrustworthy, as when, for example: 1) a document appears in two or more collections with different metadata schemes (e.g., ArXiV vs. ACM, different GitHub repositories); 2) sections/paragraphs of text are copied, as in extended journal versions of computer science papers; or 3) metadata is not curated (e.g., an ArXiV entry may not be updated to reflect a subsequent publication).
3.2. Approximate Membership Queries
Rather than attempting to directly identify pairs of similar entities, some techniques instead attempt to simply determine, online, whether a new record already belongs to the existing dataset at the time of ingestion. For example, by using a special data structure such as a hashmap or Bloom filter (Bloom, 1970) to query for matches against previously ingested records. While naive exact-matching with hashmaps or Bloom filters is efficient, it is not robust to the minor variations in text content that are inherent when using text data obtained from the web or parsed from various formats.
For text, it is preferable to use approximate notions of similarity. For instance, the Streaming Quotient Filter (SQF) (Dutta et al., 2013) is an approximate duplicate detection approach for streams of incoming data. However, the SQF is optimized for approximate duplicate detection in low-dimensional spaces and is not appropriate for text deduplication. For text, techniques such as MinHashing (Broder et al., 1998) and MinHashLSH (Leskovec et al., 2014) are more appropriate due to their ability to compress high-dimensional text records into fixed-length representations. The use of LSH functions in MinHashLSH has the additional advantage of identifying candidate duplicates using hashing as opposed to running comparisons with each existing record in the dataset. MinHash-based techniques use the Jaccard similarity to measure text overlap; however, n-gram overlap is also a reasonable metric—albeit one that is sensitive to the frequency of n-grams in a given document. Frequently, techniques evaluate the n-gram overlap of a document with the n-gram content of the remainder of the corpus using a Bloom filter, trading some granularity of comparison for space efficiency and increased throughput. This approach was utilized in preparing the DOLMA (Soldaini and others, 2024) dataset and the DataComp-LM (Li et al., 2024) baseline dataset. LSHBloom falls in this category of techniques for online approximate membership queries.
3.3. State-of-The-Art Deduplication Techniques for Text
Below, we describe in detail popular techniques for identifying duplicates in text datasets, commonly applied in real-world text dataset preparation.
@@ 3.3. State-of-The-Art Deduplication Techniques for Text
CCNet (Wenzek et al., 2020) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content. CCNet was employed as part of the data processing workflow for producing the Llama series of models (Touvron et al., 2023; Touvron and others, 2023; Dubey and others, 2024). CCNet performs dataset sharding, deduplication, language identification, and quality filtering. CCNet’s deduplication process begins by preprocessing text so that all characters are lowercase and eliminating special Unicode characters. CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates. CCNet is only capable of identifying documents that are exactly identical to one another.
@@
Dolma (Soldaini and others, 2024) deduplicates paragraphs by exact matching using a Bloom filter. The Dolma-Ngram alternative splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter. If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate. By default, $T$ = 1.0 in their configuration. These two techniques were used to produce the Dolma dataset (Soldaini and others, 2024) used to produce the Olmo series of language models (Groeneveld et al., 2024; OLMo et al., 2024).
@@
DataComp-LM (Li et al., 2024) (DCLM) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter. This technique was applied to create the DataComp-LM baseline dataset. DCLM splits each document into paragraphs on newline characters and tokenizes paragraphs with the UniSeg tokenizer (Ye et al., 2023). For each paragraph, it queries each n-gram against its Bloom filter and removes duplicated n-grams. If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document. If the proportion of duplicated n-grams in the entire document exceeds the user-defined threshold, then the document is removed from the dataset. They employ the same threshold parameters for paragraph-level and document-level deduplication.
MinHashLSH (Leskovec et al., 2014) the MinHashLSH method as described in Section 2.3. MinHash-based deduplication was applied in the data ingestion pipeline for OpenAI’s GPT-3 (Brown et al., 2020), DeepMind’s Gopher (Rae et al., 2021), AllenAI’s Olmo3 (Olmo et al., 2025), and Meta’s Llama3 (Dubey and others, 2024). MinHash-based deduplication and MinHashLSH were also employed in creating many popular language modeling datasets, such as RedPajama (Computer, 2023), the Pile (Gao et al., 2020) (used for GPT-J (Mireshghallah et al., 2023), GPT-NeoX (Black et al., 2022), and Pythia (Biderman et al., 2023)), FineWeb2 (Penedo et al., 2025), RefinedWeb (Penedo et al., 2023) (used for the Falcon series of models), StarCoder (Li et al., 2023), and StarCoder2 (Lozhkov et al., 2024). The sheer number of LLM data pipelines that employ MinHashLSH is a testament to its effectiveness. Furthermore, it highlights the potential impact of enhancing its scalability for internet-scale datasets through innovative techniques, such as LSHBloom.
4. Our Method: LSHBloom
@@
We propose a memory-efficient alternative to MinHashLSH for document deduplication. Rather than employ the traditional prefix-tree or hashmap index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in duplicate searches. The size of each Bloom filter and the number of hash functions utilized are chosen according to the method determined by Bender et al. (2022), which guarantees optimal (minimal) space usage. LSHBloom effectively acts as a Similarity-Aware Approximate Membership Query data structure, combining the compressive properties of MinHashLSH with the probabilistic indexing of Bloom filters to create a high throughput, space-efficient, online deduplication method for text ingestion workflows.
@@ 4.1. Insertion
Given a Jaccard similarity threshold $T\in$ [0, 1], MinHashLSH traditionally groups signatures into a series of $b$ bands of size $r$ such that an exact match in one band for two columns of the MinHash signature matrix denotes a duplicate document. That is, with high probability, the two documents with a matching band have Jaccard similarity greater than or equal to the threshold. To avoid storing the signatures in full, as is required by MinHashLSH, we assign each band of the minhash signature matrix its own Bloom filter. Thus, we have in total $b$ Bloom filters, and a collision in any one indicates a duplicate document.
To insert a document into the index, we first compute its MinHash signature, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (3rd and Jones, 2001)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values. Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (Carter and Wegman, 1977):
@@ 4.2. Querying
Querying a document against the index is a similar process to insertion. We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in Section 4.1. If there is a collision for any band of signatures, we mark that document as a duplicate. This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters. However, the false positive probability is configurable and can be set to be arbitrarily low. An additional benefit of LSHBloom is that the false positive overhead is analytically boundable, as we demonstrate below.
@@ 4.3. Error Rate
In traditional MinHashLSH, a collision in any band causes a document to be marked as a candidate duplicate. This is also the case in LSHBloom, where the use of Bloom filters incurs additional false positive overhead. Here, we analyze the error rate of LSHBloom. We define a false positive as a non-duplicate document that is incorrectly identified as a duplicate by LSHBloom; a false negative is a duplicate document that is not marked as a duplicate by LSHBloom.
Given a Minhash signature matrix with $b$ bands with $r$ rows each corresponding to $b$ Bloom filters with their individual false positive rates set to $p$, the net false positive overhead of using Bloom filters will be $p\textsubscript{effective}=1-(1-p)^{b}$. This is because a single collision in any band constitutes a duplicate, and false positive rates are independent for each Bloom filter. Therefore, the probability that no Bloom filter reports a false positive is $(1-p)^{b}$, and the overall false positive rate overhead introduced by using Bloom filters is the complement of this value. This means we can enforce a false positive overhead of $p\textsubscript{effective}$ by setting the value of $p$ for each Bloom filter as $p=1-(1-p\textsubscript{effective})^{1/b}$
The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows. Leskovec et al. (2014) determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
@@
Zhu et al. (2016) describe how to tune $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
@@
A Bloom filter will not yield false negatives, but may yield false positives. If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result. If MinHashLSH does not yield a false positive, the Bloom filter index may still yield a false positive result with probability ppeffective. Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.1) for insertion into the Bloom filter. Therefore, the overall false positive rate for LSHBloom is given by:
(3) $FP_{\textrm{bloom}}=FP_{lsh}+(1-FP_{lsh})*(p_{\textrm{effective}}+b/N)$
@@
Note that $p\textsubscript{effective}$ can be set arbitrarily small in practice so that this overhead is negligible. For 32-bit hashvalues (the default in packages such as datasketch) $N=2^{32}$, so the collision probability for bands before insertion into the Bloom filter is also small.
@@
Meanwhile, for each false negative that MinHashLSH produces, $p\textsubscript{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the Bloom filter index, and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion. That is,
(4) $FN_{\textrm{bloom}}=(1-(p_{\textrm{effective}}+b/N))*FN_{lsh}$
4.4. Optimizations
We codesign our implementation of LSHBloom with supercomputer hardware to maximize gains in throughput. We describe these optimizations in detail below.
4.4.1. Hashing
After profiling the original implementation of LSHBloom, we optimized the hashing operations used in LSHBloom with careful hardware and software co-design. Our profiling revealed that hashing of the integer vectors described in Section 4.1 accounted for over 90% of the time spent in the insert and query routines for LSHBloom. The key reason for this was the use of an inefficient software extended-precision integer representation included in Python, which stores extended integers as base-10 strings (Python Software Foundation, 2024). By contrast, NumPy offers fixed-precision, C-style integer representations as well as support for vectorized additions, which enhance throughput (Harris et al., 2020; NumPy Developers, 2024). However, adding hundreds of fixed-precision integers may lead to overflows, inflating the collision probability of hashvalues before insertion into our Bloom filters. The hash of each MinHash permutation is 64 bits, so adding them requires at most 71 bits of unsigned precision to perform without overflow. With this in mind, 128-bit arithmetic can be used to avoid integer overflow. Modern Intel and AMD x86_64 CPUs implement 128-bit integer arithmetic using the adc (Guide, 2011) (”add with carry“) instruction with equivalents such as addc.u64 on Nvidia PTX (Nickolls et al., 2008; NVIDIA Corporation, 2024) on other platforms. Additionally, the modular division operation performed as the last step can be implemented either using a simple bit mask or by simply using the lower order half of the 128-bit register resulting from the add with carry operation. Using Rust’s standard extended precision support (Matsakis and Klock, 2014; The Rust Project Developers, 2024) makes expressing this operation easy and results in only 30 lines of optimized x86_64 assembly depending on architecture and loop unrolling decisions. We replace our Python logic for the integer vector hashing operation described in Section 4.1 with this Rust function. This single optimized function is over $94\%$ faster than the existing Python implementation, resulting in a $11\times$ end-to-end wall clock improvement to LSHBloom.
4.4.2. Parallelization and Shared Memory
Parallelization is an obvious strategy for acceleration. As shown in Figure 1, MinHashing accounts for the majority of the runtime in LSHBloom. Thankfully, each document can be MinHashed independently and this step can easily be parallelized across multiple nodes using frameworks such as Parsl (Babuji et al., 2019). On single-node runs (as in our evaluation), we can employ multiprocessing to speed up this step. When finally writing to the Bloom filter index, we do so sequentially (without parallelism). This is because, at the time of insertion, we must also ensure that the current document is not a duplicate of a previously inserted one. Naively parallelizing this step would lead to errors in duplicate identification. Parallelizing regardless and insisting upon synchronizing Bloom filters across threads or nodes would introduce significant communication overhead. Despite this lack of parallelism, our optimized hashing routines still enable extremely fast insert and query operations. Furthermore, we take advantage of the tiny size of our Bloom filter indices by hosting our Bloom filters in node-local shared memory segments (via /dev/shm), allowing us to locate our index in DRAM with swap partitions on local SSDs. This further accelerates insert and query operations on the Bloom filter index.
@@ 4.5. Advantages
LSHBloom allows for the space-efficient deduplication of extremely large corpora. This is due to the Bloom filter’s ability to determine the set membership properties of objects (e.g., documents) of arbitrary size with a fixed number of bits. The size of the LSHBloom index only depends on the number of planned documents, $n$, the number of bands, $b$, in the MinHash signature matrix, and the desired false positive rate bound, $p$, for each Bloom filter. If we expect $n$ documents and desire a false positive bound of $p$, then each of our $b$ Bloom filters requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (Bender et al., 2022).
This means that the size of our index is invariant to the size of MinHash hashvalues—we can increase the size of MinHash hashvalues (to say, a 64 or 128-bit integer) to enhance the accuracy of our Jaccard similarity estimate at no additional cost. Meanwhile, because the traditional MinHashLSH index must store these hashvalues for comparison, increasing the size of the MinHash hashvalues results in a linear increase in the size of the traditional LSHIndex. Enhancing the accuracy of our probabilistic estimate of the Jaccard similarity by increasing the number of MinHash permutations we apply will increase the size of the MinHash signature matrix, which will only change the size of the Bloom filter index if the number of bands, $b$, in the signature matrix increases. Luckily, the number of bands in the signature matrix is generally far fewer than the number of total permutations.
@@
As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing. For these parameter settings, MinHashLSH creates nine bands (using the method described in (Zhu et al., 2016) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index. For these settings, setting a conservative ppeffective = 1e-10 for a dataset of 10B documents would only require 590 GB of space. Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex.
@@
While the use of Bloom filters to satisfy memory and throughput constraints introduces additional false positive errors, the rate of false positives and false negatives is analytically boundable. Moreover, the additional false positive error overhead scales with the number of documents we intend to insert and scales inversely with the size of each filter. Therefore, there is a trade-off between space efficiency and correctness that is mediated by the size of each filter. However, as demonstrated by our example above, this trade-off tends to be highly favorable even on extremely large datasets. Additionally, LSHBloom only introduces a single additional hyperparameter, the false positive rate for our Bloom filters, meaning that our hyperparameter tuning burden is comparable to that of MinHashLSH.
Beyond storage savings, the Bloom filter index yields substantial enhancements in ingestion throughput due to two primary factors. 1) Computational efficiency: insertion and querying against the index requires only a few independent hash operations and modular arithmetic, which is fast on modern hardware (see Section 4.4), and 2) Cache-Aware Data layout: in contrast to the pointer chasing inherent in tree or chained-hashmap-based indices, LSHBloom’s use of contiguously allocated bitarrays maximizes spatial locality and minimizes cache misses.